# OCR Hinders RAG: Evaluating the Cascading Impact of OCR on Retrieval-Augmented Generation

### v1

@@ 5 Conclusion
In this paper, we present OHRBench to evaluate the impact of OCR on RAG systems, which encompasses diverse unstructured PDF documents from six RAG applications scenarios along with Q&A pairs derived from multimodal elements in these documents. Through comprehensive evaluations of current OCR solutions, we reveal that none is fully capable of RAG systems across all scenarios. Furthermore, our analysis of different types of OCR noise demonstrates that while no retrievers and LLMs are immune to Semantic Noise, more advanced models exhibit greater resilience to Formatting Noise. We also discuss the employment of VLMs to avoid the requirement of OCR and its potential to benefit from combining image and OCR results as inputs. We believe that our document and Q&A datasets as well as perturbed data with varying levels of noise will advance future the development of OCR solutions tailored for RAG and OCR noise-resistant RAG systems.
@@ I Instruction Prompts
RAG Generation Stage Prompt Template. The prompt template for LLMs and VLMs with text-only input is shown in Tab. S4. Templates for image-only input and combined image and OCR text input are provided in Tab. S5 and Tab. S6, respectively.
Vision-Language Models OCR Prompt Template. We tune the prompt for best performance of VLMs OCR, by comparing simple and detailed instructions as shown in Tab. S7. Results in Tab. S9 indicate that the detailed prompt consistently performs better across all evaluations, so it is used by default.
@@ II.1 Document details
We collect 350 PDFs with a total of 4012 pages, of which 1370 pages are used for Q&As generation and the rest as part of the knowledge base. Each document is manually reviewed by primary authors to ensure its availability for academic use. Detailed domain statistics are shown in Tab. S1
@@
Law 60 918 505
Finance 12 1733 235
Textbook 133 133 133
Manual 15 932 225
Newspaper 116 116 116
Academia 10 204 156
Total 350 4012 1370
@@
II.2 Document with complex layout
Although existing RAG document benchmarks have gathered PDFs from different domains [14, 11, 6], they often ignore the challenges posed by OCR. Therefore, we compile PDFs with different challenging attributes in this benchmark. Specifically, we categorize these into five attributes: Multi-Elements Pages, Hard Reading Order, Complex Formula, Complex Table, and High Text Density. Detailed descriptions of each category are provided as in Tab. S2.
Attribute Counts Description
Multi-Elements Pages 1848 Pages containing at least two types of elements, such as text, images, tables, and formulas.
Hard Reading Order 1025 Pages not presented in a single-column format.
Complex Formula 1924 Pages with formulas requiring at least 10 tokens for description.
Complex Table 1283 Pages with tables having more than 4 rows or columns, or containing merged cells.
High Text Density 1003 Pages with more than 800 tokens (separated by spaces).
Table S2: Counts and description of each attribute which is present in our datasets
II.3 Ground truth structured data annotation
@@
We annotate the ground truth structured data in Mathpix Markdown format, where tables and formulas are in LaTeX and images are excluded. To ensure the quality of annotation, we begin with using Mathpix to pre-annotate all PDFs. We then remove non-semantic formatting commands to ensure unbiased ground truth for all OCR solutions. Finally, the primary authors employ Mathpix Markdown previews777 https://github.com/Mathpix/vscode-mathpix-markdown to render structured data into PDFs, manually review and correct pre-annotated results.
II.4 Rules for Formatting Noise introduction
@@
To introduce Formatting Noise, we define a perturbation rate $r$ to control its extent. In order to match the level of Semantic Noise (measured by similar edit distance), we set the $r=\{0.1,0.3,0.6\}$, indicating the three levels of perturbation: mild, moderate, and severe. Based on the Formatting Noise in the existing OCR results, we formulate the following rules to perturb plain text, tables, and formulas respectively.
II.4.1 Plain text
@@ II.4.3 Table
II.5 Rules for Semantic Noise introduction
To introduce Semantic Noise, we use samples from existing OCR results. For plain text and formula, samples are taken from MinerU, Nougat, and InternVL2-Llama3-76B. For the table, samples are from MinerU version 0.7.0b1999https://github.com/opendatalab/MinerU/tree/magic_pdf-0.7.0b1-released. To ensure diverse perturbation results, we collect 5 groups of samples and randomly select one each time. The prompt template in Tab. S8 is used to generate perturbed structured data with OCR noise. Fig. S4 presents a case.
III Additional Experimental Results
III.1 Experimental details
@@
III.2 Details in different domains
Tabs. S10, S11 and S12 present the performance of OCR solutions in retrieval, generation, and overall RAG systems across domains. The results indicate that no single OCR solution consistently outperforms others in all domains.
Figs. S1, S2 and S3 illustrate the impact of OCR noise across domains. Notably, in the finance domain, which contains many tables, as mild perturbation does not affect tables, the performance improve slightly.
@@ III.2 Details in different domains
 Given the following document, please generate three RAG-style question-answer pairs based on the document with different levels of difficulty: Easy, Medium, and Hard.
@@
 1. The question must contain all information and context/background necessary to answer without the document!!! Do not include phrases like "according to the document" in the question!
 2. The question must not contain any ambiguous references, such as ’he’, ’she’, ’it’, ’the report’, ’the paper’, and ’the document’! You MUST use their complete names!
 In your output, include the phrase from the document that contains the answer to the question as ’context’. This phrase MUST be copied verbatim, word for word, from the document. You must produce the context phrase exactly from the text, with no modifications or truncations.
@@
 - The answer form should be as diverse as possible, including [Yes/No, Numeric, Formula, Short Answer].
 - The context sentence, which is the evidence source for the question, should be as diverse as possible, including [text, table, formula].
@@
 Example Input Format: <Begin Document>…<End Document>
 Example Response: "Question": "Who was the commanding general of the Union Army during the American Civil War?","Answer": "Ulysses S. Grant", "Context": "As commanding general, Ulysses S. Grant led the Union Army to victory in the American Civil War in 1865.","Difficulty Level": "Easy" , "Answer Form": "Short Answer", "Evidence Source": "text"
User:
 <Begin Document>{document}<End Document>
@@
System:
 You are an expert, you have been provided with a question and a document image retrieved based on that question. Your task is to answer the question using the content from the given document image.
 Write the answers within <response></response>.
User:
 Question: {question}
Table S5: Prompt template for VLM generation with image-only inputs
System:
 You are an expert, you have been provided with a question, a document image, and its OCR result retrieved based on that question. Your task is to search for the content and answer these questions using both the retrieved information and the document image.
 Write the answers within <response></response>.
User:
 Question: {question}
 Retrieved Documents: {retrieved_documents}
Table S6: Prompt template for VLM generation with image+text inputs
@@
 Please do OCR on the image and give all the text content in markdown format. The formula should be wrapped in \$\$ and the table should be parsed in LaTeX format. Only output the OCR results without any extra explanations or comments.
@@
 You are a powerful OCR assistant tasked with converting PDF images to Markdown format. You MUST obey the following criteria:
@@ 1. Plain text processing:
 - Precisely recognize all text in the PDF image without making assumptions in Markdown format.
@@ 2. Formula Processing:
 - Use \$ \$ to wrap inline formulas.
 - Use \$\$ \$\$ to wrap block formulas.

### v2

@@ 5 Conclusion
In this paper, we present OHRBench to evaluate the impact of OCR on RAG systems, which encompasses diverse PDF documents from sven RAG application scenarios along with Q&A pairs derived from multimodal elements in these documents. Through comprehensive evaluations of current OCR solutions, we reveal that none is fully capable of RAG systems across all scenarios. Furthermore, our analysis of different types of OCR noise demonstrates that while no retrievers and LLMs are immune to Semantic Noise, more advanced models exhibit greater resilience to Formatting Noise. We believe that our documents featuring challenging OCR attributes and Q&A pairs sourced from varied document elements, will advance the development of OCR solutions tailored for RAG and OCR noise-resistant RAG systems.
@@ I Instruction Prompts
RAG Generation Stage Prompt Template. The prompt template for LLMs and VLMs with text-only input is shown in Tab. S9.
Vision-Language Models OCR Prompt Template. We tune the prompt for the best performance of VLMs OCR, by comparing simple and detailed instructions as shown in Tab. S10. Results in Tab. S11 indicate that the detailed prompt consistently performs better across all evaluations, so it is used by default.
@@ II.1 Document details
We curate a dataset of 1,261 PDFs spanning 8,561 pages, with 3,596 pages designated for Q&A generation and the remainder forming part of the knowledge base. These PDFs are sourced from DUDE [40], OmniDocBench [31], FinanceBench [20], CUAD [16], GNHK [22], and public resources, including Arxiv [^4] https://arxiv.org, ManualsLib [^5] https://www.manualslib.com/, LibreTexts [^6] https://libretexts.org/.
DUDE: We extract documents from the validation and test splits of DUDE, applying manual screening based on our criteria Fig. 1 to exclude samples infeasible for structured data parsing and classify each of them into 7 domains. We finally selected 450 PDFs with 4,058 images from 2,069 PDF candidates.
OmnidocBench: OmniDocBench [31] features span-level annotations and presents challenges for OCR due to its multilingual, high-resolution with dense text and handwritten content. We select all newspaper documents and manually review textbook-related samples, eliminating those with low knowledge density that hinder meaningful Q&A generation. This process yields 289 PDFs.
FinanceBench: Following prior observations [29], dboth DUDE [40] and FinanceBench [20] contain diverse document types. From FinanceBench, we randomly sample 10 PDFs characterized by large, complex tables and charts.
CUAD: We randomly select 65 PDFs to supplement the documents in law domains, which all have high text density.
GNHK: GNHK consists of handwritten documents. We manually assess and remove those with low knowledge density, finalizing a selection of 172 PDFs.
Each document is manually reviewed by primary authors to ensure its availability for academic use. Detailed domain statistics are shown in Tab. S1
@@
Law 95 1187 1143
Finance 65 2133 1359
Textbook 504 678 1126
Manual 87 1724 1155
Newspaper 279 487 1202
Academic 85 1011 1181
Administration 146 1341 1332
Total 1261 8561 8498
@@
II.2 Ground truth structured data annotation
@@
We annotate the ground truth structured data using Mathpix Markdown format, where tables and formulas are represented in LaTeX. Chart data is extracted in LaTeX table format, with charts lacking clear numeric values in figure filtered out. For images in documents, any parsable text is retained as plain text in the corresponding section. To ensure high-quality annotations, we first use Mathpix to pre-annotate all PDFs. Finally, the primary authors employ Mathpix Markdown previews777 https://github.com/Mathpix/vscode-mathpix-markdown to render structured data into PDFs, manually review and correct pre-annotated results.
II.3 Document with challenging attributes
Although existing RAG document benchmarks have gathered PDFs from different domains [18, 14, 8], they often ignore the challenges posed by OCR. To address this, we construct a benchmark that explicitly incorporates documents with challenging attributes. We define nine key attributes: structured data (tables, formulas, charts), complex layouts, handwritten content, distortions, scanned PDFs, dense text, and multilingual content. Structured data, dense text (exceeding 770 tokens), and multilingual pages are classified based on the annotated ground truth structured data. A document is considered to have a complex layout if its layout detection yields more than 20 bounding boxes. Distorted, scanned, and handwritten documents are identified during manual checks.
II.4 Q&A generation
To generate high-quality Q&A pairs covering diverse tasks and evidence sources, we define multiple prompts for each task, as detailed Tabs. S6, S7 and S8. For Chinese questions, we provide the same set of templates in Chinese to ensure that the model generates Chinese responses. Q&A with different evidence sources. For Q&A generation with evidence sourced from plaint text, table, formula and chart, we extract relevant pages from the ground truth structured data and use GPT-4o to generate Q&A pairs grounded in the corresponding evidence via tailored prompts. For Q&A related to reading order, we leverage MinerU [41], the leading model for reading order recognition [31], to identify the reading order and bounding box of paragraphs in each document. When working with documents from OmniDocBench [31], we directly use the ground truth reading order from its annotations. We verify the layout detection and reading order predictions, selecting paragraph pairs that meet one of the following criteria:
  • Adjacent paragraphs in reading order whose bounding boxes are not vertically aligned.
  • Paragraphs separated by multimodal document elements (e.g., block formulas, tables, or images).
We then randomly sample 1,500 candidate matches, manually correcting approximately 20% where MinerU’s predictions are inaccurate. We then prompt GPT-4o to generate Q&A pairs using the prompts in Tab. S7. We find that this simple prompting-based strategy can effectively generate questions with diverse evidence sources, with over 90% correctly aligned with their evidence source in our Q&A verification process.
Q&A with different tasks. To generate both understanding and reasoning questions, we apply the corresponding prompts from Tab. S7. For multi-page Q&A generation, we employ two different approaches to generate Q&A candidates: (1) Combine questions from two single-page Q&As that mention the same entity. (2) Generating multi-page questions from two paragraphs on different pages that reference the same entity. Specifically, we use spaCy [17] for named entity recognition in both single-page Q&As and document paragraphs. We then filter out candidate pairs, including: (1) Single-page Q&A pairs where the entity in one answer appears in another question. (2) Paragraph pairs that share the same entity. We finally utilize the prompts in Tab. S8 to generate multi-page questions. However, despite the many optimizations of the prompt and generation strategies we tried, GPT-4o sometimes produces Q&A pairs that are either answerable with a single paragraph or simply concatenate two single-page questions while maintaining separate evidence sources instead of high-quality and realistic multi-page Q&As. To address these limitations, we develop a comprehensive filtering process to ensure the quality of multi-page Q&As, as detailed in Sec. II.5.
II.5 Q&A verification.
We verify Q&A quality based on three criteria: (1) Compatibility with realistic RAG applications, (2) faithfulness to task definition, and (3) correctness. Below, we detail our approach for each aspect.
Compatibility with Realistic RAG Applications. To assess context dependence, we identify key patterns from existing context-dependent questions and apply the following heuristics:
  • Questions lacking an explicit entity name.
  • Questions containing more than one ambiguous pronouns (e.g., "he," "she," "it," "they", "this", "that").
  • Questions featuring phrases such as "in the document" or "according to the document."
These rules filter most context-dependent questions. We then refine the selection using prompts in VisRAG [47] and DeepSeek-V3 to further distinguish context-dependent questions from the remaining set. Additionally, we use GPT-4o to exclude questions answerable without retrieval by instructing it to respond without providing evidence context across both single-page and multi-page Q&As.
Faithfulness to Task Definition. Based on the Q&A verification prompts in [10], we use the prompts in Tab. S12 to assess faithfulness using DeepSeek-V3. To verify the validity of evidence sources, we locate them in the original ground truth structured data and ensure they originate from the correct corresponding LaTeX code environments. For the multi-page and reading-order questions, we employ GPT-4o to generate three responses: (1) without context, (2) with context A, and (3) with context B. If any response yields a correct answer, the question is excluded, ensuring that only truly multi-page or reading-order-related questions remain.
Correctness. To guarantee each Q&A has a unique and correct answer supported by its evidence context, we provide oracle evidence and sample GPT-4o’s response 10 times. We apply a best-of-N strategy to determine the final answer, which must match the ground truth. Q&As with fewer than three consistent correct responses are also excluded.
Our filtering pipeline underwent two iterations of refinement. In each round, we randomly sample 100 Q&As to verify the filtering results adherence to our criteria. Finally, to mitigate false positives, we manually reviewed all remaining questions, yielding 8,498 high-quality Q&As from an initial pool of 15,317 candidates.
III OCR Noise Introduction
III.1 Rules for Formatting Noise introduction
@@
To introduce Formatting Noise, we define a perturbation rate $r$ to control its extent. In order to match the level of Semantic Noise (measured by similar edit distance), we set the $r=\{0.1,0.3,0.6\}$, indicating the three levels of perturbation: mild, moderate, and severe. Based on the Formatting Noise in the existing OCR results, we formulate the following rules to perturb plain text, tables, and formulas, respectively.
III.1.1 Plain text
@@ III.1.3 Table
OCR Avg. Counts
MinerU 35.0
GOT 45.7
Nougat 63.2
F-minor 37.9
F-moderate 42.2
F-severe 56.3
Table S2: Counts of Formatting Noise. The counts of Formatting Noise we add (F-minor, F-moderate, F-severe) is approximately the distribution of the counts of Formatting Noise for MinerU, GOT and Nougat.
III.2 Rules for Semantic Noise introduction
In order to construct perturbed document images that conform to the realistic distribution of naturally distorted documents, we use a cross-validated process involving multiple annotators. We finally identify 8 strategies from [3] as follows:
  • Background Addition: We collect 15 background images of real paper textures and blend them with original images at an 80:20 ratio.
  • Salt-and-Pepper Noise: Randomly replace 1% of the image pixels with white ("salt" noise) and black ("pepper") pixels.
  • Dirty Rollers: Add random rollers with thickness between 1 and 3 pixels, a line addition probability of 0.05 per pixel row or column.
  • Random Rotation: Apply a random rotation of $-3^{\circ}$ and $+3^{\circ}$.
  • Binarization: Utilize the Augraphy [^9] https://github.com/sparkfish/augraphy to simulate effects such as dilation, erosion, and letterpress printing.
  • Warping: Apply geometric transformations and folding effects via Augraphy to mimic paper creases.
  • Shadows: Apply light gradients and shadow cast from Augraphy to simulate shadows that occur when a document is taken.
  • Blur via Point Spread Function: Generated 100 PSF kernels and randomly applied one to the document.
We classify above distortions into two categories: (1) weak distortions: These preserve text clarity and include background addition, binarization, minor rotation, and PSF-based blurring. (2) strong distortions: These degrade readability, causing blurriness and font warping. They include salt-and-pepper noise, dirty rollers, warping, and shadow effects. To simulate varying levels of document distortion, we apply the above strategies in three ways:
  • Apply a weak distortion per page.
  • Apply a strong distortion per page.
  • Apply two randomly selected distortions per page.
We generate three document image datasets with varying noise levels and parse structured data using MinerU, GOT, and Qwen2.5-VL, resulting in 9 perturbed datasets. The examples of distorted documents are shown in Fig. S1. In Sec. 4.3, we evaluate RAG performance on these datasets, reporting the average results for each noise level.
IV Additional Experimental Results
IV.1 Experimental details
@@
IV.2 Details in different domains
Domain GT MinerU Marker GOT Nougat Qwen2.5-VL InternVL2.5
Law 81.2 71.0 77.1 62.1 69.0 76.4 69.6
Finance 59.7 36.4 45.0 30.4 25.8 47.9 47.1
Textbook 73.2 43.8 49.6 48.8 37.1 58.3 55.0
Manual 79.1 60.4 68.6 58.9 47.8 71.3 70.2
Newspaper 40.5 31.3 34.0 12.4 10.6 27.7 18.4
Academic 75.1 50.3 55.2 50.2 45.0 61.1 57.1
Administration 82.2 59.4 68.3 57.7 52.7 73.1 73.8
All 70.0 50.1 56.6 45.4 40.8 59.2 55.8
Table S3: Retrieval performance across different domains.
Domain GT MinerU Marker GOT Nougat Qwen2.5-VL InternVL2.5
Law 56.9 53.4 54.4 43.3 48.8 53.9 50.9
Finance 43.1 30.1 29.5 19.7 17.7 35.9 36.8
Textbook 37.6 25.9 28.2 24.8 16.8 29.1 29.1
Manual 50.2 45.3 46.1 41.3 34.3 48.7 47.7
Newspaper 35.0 33.7 31.6 9.5 8.4 19.6 11.7
Academic 38.3 29.5 27.9 25.3 24.8 33.2 31.3
Administration 46.4 35.7 37.7 32.2 29.2 42.7 42.9
All 43.9 36.1 36.3 27.8 25.5 37.5 35.8
Table S4: Generation performance across different domains.
Domain GT MinerU Marker GOT Nougat Qwen2.5-VL InternVL2.5
Law 49.6 48.1 48.1 41.1 43.9 47.2 44.9
Finance 27.2 19.4 20.1 15.1 13.1 22.9 22.8
Textbook 30.5 20.9 22.5 21.0 15.7 23.8 23.5
Manual 44.4 38.1 39.8 36.0 30.7 42.3 41.6
News 29.0 25.6 24.7 8.3 5.6 17.4 11.0
Academic 31.9 25.6 24.1 22.8 21.2 27.6 26.4
Administration 41.0 30.9 32.7 29.2 26.6 37.3 37.5
All 36.1 29.5 30.0 24.6 22.2 31.1 29.6
Table S5: Overall performance across different domains.
Figure S1: Cases of distorted documents.
@@ IV.2 Details in different domains
 You are an AI specialized in generating QAs from documents. Your mission is to analyze the document, follow the instructions, and generate RAG-style question-answer pairs based on the document.
@@
 1. Question should represent a plausible inquiry that a person (who has not seen the page) might ask about the information uniquely presented on this page. The questions should not reference this specific page directly (by page number, pointing to a specific paragraph or figure, and never refer to the document using phrases like ’in the document’), nor should they quote the text verbatim. They should use natural language reflecting how someone might inquire about the page’s content without direct access.
 2. Question must contain all information and context/background necessary to answer without the document. Do not include phrases like "according to the document" in the question.
 3. Question must not contain any ambiguous references, such as ’he’, ’she’, ’it’, ’the report’, ’the paper’, and ’the document’. You MUST use their complete names.
User:
 Your task is to generate several RAG-style question-answer pairs with different levels of difficulty and evidence sources. {detailed_task_description}.
@@
 - The answer form should be as diverse as possible, including [Yes/No, Numeric, String, List].
 - {additional_task_criteria}
@@
{qa_examples}
 <document>{document}</document>
@@
Structure data task:
 In the given documents, the chart elements are all enclosed within <chart> </chart> tags and illustrated in LaTeX table format. Pay attention to the difference between them and tabular data, as tabular data is not enclosed by <chart> </chart> tags. # This paragraph is only used for chart data.
 In order to generate this type of question-answer pairs, first, you need to read the given document, identify the table/formula/chart elements within it, and use them as the evidence context. The evidence context can be a single paragraph for single-hop questions, or several related paragraphs for generating multi-hop questions that require reasoning. After that, you need to generate questions and corresponding answers based on them.
Reading order task:
 Your task is to generate RAG-style question-answer pairs from the given two documents.
In order to generate this type of question-answer pairs, first, you need to read the given two documents (A, B), identify the text sharing the same entities, and design a question-answer pair based on the contents of both documents A and B. If it is based on the message of document A or document B alone, it cannot be answered.
Understanding task:
 You should generate question-answering pairs that require the responser to extract information from documents. The answer should be able to find directly in the documents without any reasoning.
Reasoning task:
 You should generate question-answering pairs that require responser to reason before answering, such as calculations, comparisons, finding the maximum and minimum, or integration information from different parts of the documents. The answer should not be able to be found directly in the documents.
Table S7: Detailed description used to generate Q&A pairs for different tasks.
Multi-page Q&A from single-page question:
 Your mission is to generate RAG-style combined questions from two questions that have the same entity.
 When generating a combined question, there are some criteria you should follow:
 - The answer to the combined question should be the same as the answer2.
 - It must combine the answer1 to question1 to answer the combined questions. This means that, to answer the combined question, a responder must first deduce the part of the combined question that refers to the answer1, and then proceed to answer the combined question based on that answer.
 - You cannot include the answer to question 1 in the combined question.
{combined_qa_examples}
 Based on the above 3 examples, provide a combined question for the following case. If you find it is hard to create such a combined question, output None as the answer. Enclose the combined question within <answer> </answer>:
 question1: {q1}
 answer1: {a1}
 question2: {q2}
 answer2: {a2}
Multi-page Q&A from different paragraphs:
 Your task is to generate RAG-style question-answer pairs from the given two documents and entity names. The entity names appear in both documents, and you need to use them as a bridge to generate the RAG-style question-answer pairs that need to be answered by combining information from both documents.
 To generate the question-answer pairs, first, you need to read the given two documents (A, B) and the entity names, find paragraphs related to them, use the paragraphs as evidence context, and design a question-answer pair based on the evidence context from the two documents.
Table S8: Detailed description used to generate multi-page Q&A pairs from both single-page questions and different paragraphs sharing same entities.
@@
 Please do OCR on the image and give all the text content in markdown format. The formulas should be wrapped in \$\$. The table and charts should be parsed in LaTeX format. Only output the OCR results without any extra explanations or comments.
Table S10: Simple prompt for VLMs OCR
@@
 You are a powerful OCR assistant tasked with converting PDF images to the Markdown format. You MUST obey the following criteria:
@@ 1. Plain text processing:
 - Accurately recognize all text content in the PDF image without guessing or inferring.
 - Precisely recognize all text in the PDF image without making assumptions in the Markdown format.
 - Maintain the original document structure, including headings, paragraphs, lists, etc.
@@ 2. Formula Processing:
 - Enclose inline formulas with \$ \$. For example: This is an inline formula \$ E = mc2̂ \$.
[…]